OpenStudy (anonymous):

Just need confirmation again (2t^3)^4 + 2t^12?

OpenStudy (anonymous):

It's supposed to be a equals sign!

OpenStudy (anonymous):

Sorry.

OpenStudy (anonymous):

first you need to apply this rule of exponents to the left side of that equation: (a^m)^n=a^m*n

OpenStudy (anonymous):

if you did that, then you are correct

OpenStudy (anonymous):

If the whole thing is raised to the power of 4 as suggested by the () then that would include the 2 in the base as well.. so 2^4 is also necessary

OpenStudy (anonymous):

What...? o.o

OpenStudy (anonymous):

If the problem has parenthesis then it means each thing in the parenthesis is to use the exponent. You can think of them separately... (2t^3)^4 = [2^4][(t^3)^4]

jimthompson5910 (jim_thompson5910):

\[\large (2t^3)^4\] \[\large (2t^3)(2t^3)(2t^3)(2t^3)\] \[\large (2*2*2*2)(t^3*t^3*t^3*t^3)\] \[\large 16t^{3+3+3+3}\] \[\large 16t^{12}\] So \[\large (2t^3)^4=16t^{12}\]